(6) \theta  \epsilon   \mathbb{R}. We pick n = 2m +1m > 0) samples randomly from the population which distributes uniformly between the  interval of. 十 . Answer the following questions regarding the median of the samples \widetilde{X_{n}}

\cdot  Show that the distribution which \widetilde{X_{n}} follows has the distribution as shown below.

Lon(x) = (2m +1}{4-6 – 017]     (0 - <x<0+

\cdot   Find the expected value of \widetilde{X_{n}} .

\cdot  Show Var[xn = 4(2m +3) .

\cdot When n → , show that \widetilde{X_{n}} is the consistent estimator of \theta.
Answer #1

X2 0 - 0 + ) f(x) = 0 I 0-1 cocotte ow F(x) = x-10 X-te n. 2 m +1 cm 20) that is n is odd. So middle observation is (m+yth ob90(x) = (2)! (x+6+4)(-14.04). ml ml em til mimi - Comment [(x-any)(1-346)] (2m to! mim! Id= 1b-x-a we have which is (athyExpectation E (Ăn) = Q+ of Rn ( x. 9(4). Note that in ca) also written as, ancap (mel,m4u) (x-04L) [1-(4-044) 7 at Ecznat sEXO) - a s BIMH, MAD) ymil(1-) dy + (0-4) tym clcym dy form #21m +1) +(0-) pima, mte) BIMAI, MAIL BIM +2,m11) Bimti, m+0) 0-ECTn) - $10 Ecão) = 0 Son is unbiased estimator of O. | Vd (X) = (xn) - [E (Zn) EG) - Ix? gn (x) dx (x-ot! BCmtim using subsCan? I s pemil, mi) BCM 13, mai) +20-1) B (1942, tad 460-) p (mtile rot1) 7 BCM + 3, m+1) 11201) BM 42 inti), PIMAI, mi) BermVarcxn) – ECGN)? (ECEN)? Mt2 2 12m43) to 2 + (0-1)-02 Mt2 2 (RM43) + 0.4+0244-0-02 Mt2 202 M +3) hmt 8-hm-6 8 (2m 13) 8t2m+3)